# Faces of Fairness: Examining Bias in Facial Expression Recognition Datasets and Models

### v1

@@ 4.1.5 Dataset Generality
To investigate general bias in FER datasets, we conducted an experiment where a model was trained to identify the dataset to which any given image belongs. Each image was labeled with its respective dataset name, and the XceptionNet model [43] was trained to predict the dataset label. The results, as shown in Fig. 6, reveal biases inherent to FER datasets. While some correlation exists between RAF-DB and two other datsets (ExpW and AffectNet), Fer2013, ExpW, and AffectNet remain uncorrelated. This experiment underscores that differences in data collection methodologies—including image acquisition, processing, cropping, and storage—significantly affect FER datasets. Such biases, stemming from dataset-specific protocols, remain a relatively underexplored drawback in FER research.
TABLE II: Bias analysis across different datasets (in percent). When one dataset is excluded, a model is trained on the remaining datasets. Model is evaluated on the validation set of the included datasets, and then is tested over the excluded dataset. More difference between validation and test scores highlights more inherent biases in that dataset.
@@
An additional experiment was conducted to further investigate the generality and biases inherent in the datasets. In this experiment, one dataset was excluded, and a facial expression recognition model (XceptionNet) was trained using the remaining datasets. To ensure fairness, the model was first evaluated on the unseen validation sets of the datasets included in training. Subsequently, the generalizability of the excluded dataset was assessed by testing the model on its images. Since the excluded dataset was not part of the training process, the results reflect the dataset’s bias. Table II compares the validation accuracies achieved on the included datasets with the test accuracy for the excluded dataset. Both Validation and Test accuracies in the table represent average per-class accuracies. Interestingly, the Test accuracies for RAF-DB and Fer2013 were higher than their Validation accuracies, while the opposite trend was observed for the ExpW and AffectNet datasets. For instance, when AffectNet was excluded, the Validation accuracy across the validation sets of RAF-DB, Fer2013, and ExpW was 57.0%. However, the Test accuracy on the excluded AffectNet dataset dropped to 45.9%. This result suggests that the challenges presented in the AffectNet dataset are not sufficiently captured by RAF-DB, Fer2013, and ExpW. A similar pattern was observed when the ExpW dataset was excluded. Conversely, for RAF-DB and Fer2013, the Validation accuracies were lower than their respective Test accuracies, demonstrating an inverse trend. Based on these findings, we conclude that the generalizability of the AffectNet and ExpW datasets is higher than that of RAF-DB and Fer2013. This indicates that RAF-DB and Fer2013 exhibit more inherent biases compared to AffectNet and ExpW.
@@ 4.2 Fairness of Models
After analyzing bias in the FER data, we further investigate bias and fairness within the models. As mentioned, four bias metrics are used to evaluate six deep models with different architectures. In the following sections, we describe our training methodology and then discuss the accuracy and bias of each model.
Figure 7: Model accuracy for each expression (in percent). In this experiment, models were trained on individual datasets and evaluated on their respective evaluation sets. $\mu$ and $\sigma$ denote the average accuracy and its standard deviation for each dataset, while MEAN and STD represent the overall average accuracy and standard deviation across each model. RAF-DB emerged as the least challenging dataset, whereas ExpW posed the greatest difficulty. Among the models, GPT-4o-mini delivered the highest accuracy, contrasting with the CLIP model, which recorded the lowest performance.
@@ 4.2.1 Training and Implementation Details
In this experiment, we selected six deep learning architectures, including both CNN- and transformer-based models: MobileNet, ResNet, XceptionNet, ViT, CLIP, and GPT-4o-mini. MobileNet was used as a baseline CNN model due to its streamlined feed-forward structure without residual connections, making it computationally efficient for mobile and embedded applications. ResNet and XceptionNet, incorporating residual connections, were chosen for their ability to mitigate the vanishing gradient problem and improve model accuracy. Among the transformer-based models, ViT represented a state-of-the-art approach for computer vision tasks, leveraging self-attention mechanisms instead of traditional convolutional operations. CLIP, a large vision-language model designed for multi-modal learning, enabled joint processing of images and text. Lastly, GPT-4o-mini, a lightweight variant of GPT-4, was included for its capability to efficiently handle both NLP and vision tasks.
@@
In the first experiment, we fully trained four models, including MobileNet, ResNet, XceptionNet, and ViT, using each dataset and then evaluated their accuracy over their corresponding validation set. In the next experiment, we concatenated training sets, trained all the models, and evaluated them using the concatenated validation set. All four models were trained for 30 epochs with a fixed batch size of 35. The Adam optimizer was used with a learning rate of $10^{-3}$ for MobileNet, ResNet, and XceptionNet, and $10^{-5}$ for the ViT model. The two other Adam parameters were set to beta-1 = $0.9$ and beta-2 = $0.999$ for all models. Notably, we selected the categorical cross-entropy loss function for our training process.
Aside from the ViT model, we also leveraged two other transformer-based models, CLIP and GPT-4o-mini. These two models are large vision-language models capable of handling multiple tasks. Since CLIP is trained on multi-modal datasets and is open-source, we fine-tuned its decision-making head using each training set of different datasets, and their combined version. Then we evaluated its accuracy and fairness on each dataset and the concatenated version. We used the Adam optimizer with a learning rate of $10^{-3}$, beta-1 of $0.9$, and beta-2 of $0.999$. Additionally, we selected the categorical cross-entropy loss function for fine-tuning this model over 30 epochs with a batch size of 35.
GPT-4o-mini is another transformer-based vision-language model included in this research. However, since it is not open-source, we did not train or fine-tune it for this specific task. To evaluate this model and compare its accuracy and fairness with the other models, we only evaluated it using the validation set of each dataset and their combined version. We prompted all the facial images of the validation set into this model and asked it to assign one of the seven expressions used in this study. The prompt we used was: ”What is the expression of this person among happy, sad, surprise, fear, disgust, anger, and neutral, in only one word? Try to find one of these expressions, but if your prediction is out of these expressions, select the closest expression to it (among the 7 aforementioned expressions). If you cannot determine any expression, randomly select one of the mentioned expressions. If the predicted expression is another word in the family of the aforementioned expressions (happy, sad, surprise, fear, disgust, anger, neutral) change it to the corresponding word. For example, fearful is equal to fear, disgusted is equal to disgust, angry is equal to anger, surprised is equal to surprise, scared is equal to fear, and so on. If you randomly select an expression, do not need to explain, only generate the expression name. We can leverage this data for comparing their accuracy, and more importantly, models’ bias over each group of data, on over different demographic attributes, age, gender, and race.
@@ 4.2.2 Facial Expression Recognition
Figure 8: Accuracy of each model across all datasets (in percent). In this experiment, all datasets were used for both training and validation. Notably, no model was able to achieve high scores for the two challenging expressions, Fear and Disgust. A comparison of the results reveals that GPT-4o-mini and ViT models demonstrated the most consistent accuracy scores, while CLIP struggled to achieve satisfactory performance.
TABLE III: General accuracy of the models (in percent). Four models—MobileNet, ResNet, XceptionNet, and ViT—were trained on the full training samples of all datasets and evaluated using their respective validation sets. The CLIP model was fine-tuned on the training sets and validated using the validation set, while GPT-4o-mini was evaluated solely on the validation sets, without any training or fine-tuning.
@@
The average accuracy for each dataset, represented by $\mu$ in Fig. 7, demonstrates a correlation between dataset Richness and complexity. Richer datasets tend to be more challenging, leading to lower average accuracies. The last two columns of this figure, labeled Mean and STD, summarize the mean accuracy and standard deviation for each model. These results show GPT-4o-mini achieved the highest accuracy while maintaining the lowest standard deviation. This performance can be attributed to GPT-4o-mini’s extensive pretraining across diverse tasks and modalities, enabling it to extract features particularly well-suited for vision tasks. While this suggests GPT-4o-mini may be the fairest model, subsequent experiments reveal it exhibits significant biases.
In contrast to GPT-4o-mini, the CLIP model showed the worst performance, with an average accuracy below 50% and a high standard deviation exceeding 25%. These results indicate that the CLIP model is not well-optimized for FER tasks in its current form. Training the entire model, rather than just fine-tuning its decision-making head, could potentially improve its performance. Lastly, among the fully trained models (MobileNet, ResNet, XceptionNet, and ViT), ViT achieved the highest accuracy score, while ResNet exhibited the lowest standard deviation. Based solely on high accuracy and low standard deviation metrics, the models can be ranked from least to most accurate models as follows: GPT-4o-mini, ViT, ResNet, MobileNet, XceptionNet, and CLIP. We will discuss this ranking in the next sections, when we extract the metric-based bias scores.
A closer analysis reveals significant confusion among certain expressions. For instance, pairs like Sad-Neutral, Disgust-Neutral, and Disgust-Anger pose considerable challenges, with erroneous predictions in the range of [18.8-35.3] percent, while zero values would be expected. While Fig. 7 showed GPT-4o-mini as the top performer across various datasets, the general model results, summarized in Table III, indicate that the ViT model achieved the highest average accuracy in this experiment. Additionally, MobileNet had the lowest standard deviation, indicating more consistency in its predictions. For further details on the experiments, please visit our GitHub.
So far, we have assessed the models’ accuracy to evaluate their robustness against bias stemming from imbalanced data distribution. Our findings indicate that, overall, GPT-4o-mini and ViT are the most accurate models, while CLIP performed poorly. Although these results suggest that GPT-4o-mini and ViT might be the least biased models, subsequent experiments reveal different outcomes. In the next step, we will apply various bias metrics to further analyze the models’ biases across different attributes.
4.2.3 Bias Analysis
To investigate bias, we combined the training sets of all four datasets into a unified training set, applying the same approach to the validation sets. Four models—MobileNet, ResNet, XceptionNet, and ViT—were trained on the unified training set, and their bias scores were evaluated using the unified validation set. Additionally, the CLIP model was fine-tuned on the training set, and its bias scores were calculated using the validation set. For the GPT-4o-mini model, bias scores were directly quantified using the validation set without any additional training or fine-tuning. This section first explores bias across different attributes and expressions for each model, followed by a comparison of the models’ overall bias. It is notable that all the four bias metrics used Eq. 12 for their calculations.
TABLE IV: Bias analysis using Equalized Odds, in the range of [0, 100], across models and three demographic attributes: age, gender, and race. Bias is calculated for each expression, followed by the calculation of maximum value, mean, and standard deviation for each attribute.
@@ 4.2.3 Bias Analysis
TABLE V: Bias analysis using Equal Opportunity, in the range of [0, 100], across models and three demographic attributes: age, gender, and race. Bias is calculated for each expression, followed by the calculation of maximum value, mean, and standard deviation for each attribute.
@@
The first metric, Equalized Odds, examines the fairness of True Positive (TP) and False Positive (FP) rates across different groups. Table IV highlights that the Neutral and Happy expressions were the most biased, with nearly all maximum biases for age, gender, and race originating from these two expressions. For the Age attribute, Surprise emerged as the least biased expression, whereas for Gender and Race attributes, Fear showed the lowest bias. The Max column of this table reveals that the Age attribute consistently exhibited the lowest bias across all models. In contrast, Gender displayed the highest bias in four out of six models. However, for the CLIP and GPT-4o-mini models, Race was the most biased attribute. The Mean and STD columns further emphasize that, on average, Race was the most biased attribute across all models, while Age demonstrated the least bias. When comparing the Mean and STD values among the models, the Equalized Odds metric indicates that the ViT model was the most biased overall, while ResNet demonstrated the lowest bias score.
TABLE VI: Bias analysis using Demographic Parity, in the range of [0, 100], across models and three demographic attributes: age, gender, and race. Bias is calculated for each expression, followed by the calculation of maximum value, mean, and standard deviation for each attribute.
@@
TABLE VII: Bias analysis using the Treatment Equality, in the range of [0, 100], across models and three demographic attributes: age, gender, and race. Bias is calculated for each expression, followed by the calculation of maximum value, mean, and standard deviation for each attribute.
@@
Another metric we introduced in Sec.3.2 is Equal Opportunity, which focuses on the ratio of TP predictions across different groups. Information about this metric is provided in Table V. Unlike Equalized Odds, this table shows that the Happy expression was not the most biased in terms of Equal Opportunity. For the Age attribute, Neutral ranked highest in bias, while for Gender and Race attributes, Happy and Surprise expressions exhibited the highest bias scores, respectively. Examining the Max, Mean, and STD columns of this table reveals that Race is the most bias-affected attribute across all models. Based on this metric, GPT-4o-mini exhibited the highest bias, while ResNet and MobileNet achieved the lowest bias scores. Comparing the Mean values from Tables IV and V, it becomes evident that Equal Opportunity demonstrates more bias than Equalized Odds. For instance, the bias range for Age in Equalized Odds was [1.7-2.3], while Equal Opportunity showed a higher bias range of [3-4.6].
Demographic Parity analyzes the likelihood of any group receiving a particular outcome. The data in Table VI shows that, compared to other bias metrics, Demographic Parity results in lower bias scores. For the Age attribute, the highest bias was observed in the Neutral expression, while Happy, Sad, and Surprise had minimal effects on the Demographic Parity of Age. In contrast, Happy was the most biased expression for the Gender attribute, while Fear and Sad showed the least bias. For the Race attribute, Neutral and Happy were the primary sources of bias, while Fear exhibited the lowest bias across all models. Although the Max column highlights higher biases for the Gender attribute, the Mean column underscores Race as the most biased attribute overall. As with the other bias metrics, Age remained the least biased attribute. The Mean and STD columns illustrate that for the Age attribute, the CLIP model was the most biased model, while for Gender and Race, the ViT model was the most biased model. Overall, considering the Mean bias scores, the ViT model emerged as the most biased, while ResNet demonstrated the least bias.
The final bias measure we examined is Treatment Equality, which assesses the distribution of misclassified samples across different groups. Table VII demonstrates that Treatment Equality resulted in higher bias scores. A detailed expression-level analysis shows that Sad, Fear, and Disgust were the main sources of bias for the Age attribute. While Happy and Anger exhibited the highest biases for the Gender attribute, Happy significantly influenced the bias score for Race in three out of the six models. Among the attributes, Race had the highest bias scores, while Age exhibited the lowest. This trend was consistent with our findings in the Equalized Odds and Equal Opportunity metrics. Comparing the Max column in Table VII with the Max columns in Tables IVV, and VI shows that Treatment Equality exhibited higher bias scores across all attributes. In a model-wise comparison of Treatment Equality, CLIP showed the lowest bias score, while the highest bias was observed in GPT-4o-mini.
TABLE VIII: Bias score for all models using the measures Equalized Odds (Eq-Od), Equal Opportunity (Eq-Op), Demographic Parity (De-Pa), and Treatment Equality (Tr-Eq) across age, race, and gender attributes. The mean represents the average of each row, and the bias score is the final mean score of the attributes. Data are in the range of [0, 100].
@@
To sum up, we analyzed the bias of the models for each attribute and expression, followed by a model-based bias comparison. Based on our experiments, the ViT model showed the highest bias in the Equalized Odds and Demographic Parity metrics, while GPT-4o-mini exhibited significant bias in the Equal Opportunity and Treatment Equality metrics. On the other hand, ResNet was the most robust model against bias in three out of four bias metrics.
To generate a unique mathematical bias score, we utilized Eqs. 13 and 14 to calculate the final bias score for each model. Table VIII presents the bias scores of each model across different attributes and bias metrics. As outlined in Eq. 12, we used the maximum bias score of each model (i.e., the Max column from Tables IVVVI, and VII) as the bias score for each model. In the next step, we calculated the average bias score for each model across all attributes using Eq. 13 (the average of each row in Table VIII). Finally, the overall bias score for each model was obtained by averaging the bias scores for each attribute by Eq. 14. In conclusion, the results showed that GPT-4o-mini was the most biased model, followed by ViT in second place. Additionally, this table highlights that residual-based models, including ResNet and XceptionNet, were the least biased models. Ultimately, our experiments rank the models from the most biased to the least biased as follows: GPT-4o-mini, ViT, CLIP, MobileNet, XceptionNet, and ResNet. Notably, this research focused solely on the FER task, and the observed bias in the models could be evaluated on different tasks, potentially yielding varying bias scores.

### v2

@@ 4.1.6 Dataset Generality
To investigate general bias in FER datasets, we conducted an experiment where a model was trained to identify the dataset to which any given image belongs. Each image was labeled with its respective dataset name, and an XceptionNet model [30] was trained to predict the dataset label. The results, as shown in Fig. 6, demonstrated biases inherent to FER datasets. While some correlation exists between RAF-DB and two other datsets (ExpW and AffectNet), Fer2013, ExpW, and AffectNet remain uncorrelated. This experiment underscores that differences in data collection methodologies—including image acquisition, processing, cropping, and storage—significantly affect FER datasets. Such biases, stemming from dataset-specific protocols, remain a relatively underexplored drawback in FER research.
TABLE III: Bias analysis across different datasets (in percent). When one dataset is excluded, a model is trained on the remaining datasets. Model is evaluated on the validation set of the included datasets, and then is tested over the excluded dataset. More difference between validation and test scores highlights more inherent biases in trained datasets.
@@
An additional experiment was conducted to further investigate the generality and biases inherent in the datasets. In this experiment, one dataset was excluded, and a facial expression recognition model (XceptionNet) was trained using the remaining datasets. To ensure fairness, the model was first evaluated on the unseen validation sets of the datasets included in training. Subsequently, the generalizability of the excluded dataset was assessed by testing the model on its images. Since the excluded dataset was not part of the training process, the results reflect the dataset’s bias. Table III compares the validation accuracies achieved on the included datasets with the test accuracy for the excluded dataset. Interestingly, test accuracies for RAF-DB and Fer2013 were higher than their validation accuracies. This indicates that two other datasets, ExpW and AffectNet used for training, have provided good generalability. An opposite trend was observed for ExpW and AffectNet which shows that when we used RAF-DB and Fer2013 as train sets the model could not capture challenhes in ExpW and AffectNet datasets. For instance, when AffectNet was excluded, validation accuracy across the validation sets of RAF-DB, Fer2013, and ExpW was 57.0%, while test accuracy on the excluded AffectNet dataset dropped to 45.9%. This experiment gives us clues for more inherent bias over RAF-DB and Fer2013 datasets, and lower bias over ExpW and AffectNet.
4.1.7 Conclusion
The analysis of the four FER datasets reveals that all exhibit notable demographic and expression biases. None of the datasets are fully balanced across Age, Gender, Race, or expression categories, and these imbalances may affect model performance and fairness. While the degree and nature of bias vary slightly between datasets, the overall patterns are consistent, highlighting the importance of considering dataset composition when developing and evaluating facial expression recognition models.
@@ 4.2 Fairness of Models
After analyzing bias in the FER datasets, we further investigate bias and fairness within the models. As mentioned, four bias metrics are used to evaluate eight deep models with different architectures. In the following sections, we describe our training methodology and discuss the accuracy and bias of each model.
Figure 7: Model accuracy for each expression (in percent). In this experiment, models were trained on individual datasets and evaluated on their respective evaluation sets. $\mu$ and $\sigma$ denote the average accuracy and its standard deviation for each dataset, while MEAN and STD represent the overall average accuracy and standard deviation across each model. RAF-DB emerged as the least challenging dataset, whereas ExpW posed the most difficulty. Among the models, POSTER reached the highest accuracy, contrasting with CLIP which showed the lowest accuracy.
@@ 4.2.1 Training and Implementation Details
We selected eight deep learning architectures, including six basic CNN- and transformer-based models: MobileNet, ResNet, XceptionNet, ViT, CLIP, and GPT-4o-mini, plus two SOTA FER-specialized models: POSTER and CEPrompt. MobileNet was used as a baseline CNN model due to its streamlined feed-forward structure without residual connections, making it computationally efficient for mobile and embedded applications. ResNet and XceptionNet, incorporating residual connections, were chosen for their ability to mitigate the vanishing gradient problem and improve model accuracy. Among the transformer-based models, ViT represented a state-of-the-art approach for computer vision tasks, leveraging self-attention mechanisms instead of traditional convolutional operations. CLIP, a large vision-language model designed for multi-modal learning, enabled joint processing of images and text. GPT-4o-mini, a lightweight variant of GPT-4, was included for its capability to efficiently handle both NLP and vision tasks. To better understand bias in FER, we examined the POSTER model, which integrates self-attention with posterior distribution modeling to capture subtle emotion cues, and the CEPrompt model, which applies context-aware prompts to guide vision transformers for more robust, context-sensitive facial expression recognition.
@@
In the first experiment, we evaluated each model’s accuracy on every dataset separately. In the second, we merged all training sets, trained each model, and then measured accuracy on the combined validation set. Finally, the third experiment focused on assessing each model’s bias utilizing the bias metrics. For this, we fully trained four models—MobileNet, ResNet, XceptionNet, and ViT—and fine-tuned CLIP, POSTER, and CEPrompt. Full training was conducted for 30 epochs, while fine-tuning was limited to 5 epochs. We used the Adam optimizer with a learning rate of $10^{-3}$ for MobileNet, ResNet, XceptionNet, CLIP, and CEPrompt, and $10^{-5}$ for ViT and POSTER. Categorical cross-entropy served as the loss function for all training procedures.
Since GPT-4o-mini is not open source, we evaluated it without any fine-tuning. We prompted all the facial images of the validation sets into this model and asked it to assign one of the seven expressions used in this study to that image. The prompt we used was: ”What is the expression of this person among happy, sad, surprise, fear, disgust, anger, and neutral, in only one word? Try to find one of these expressions, but if your prediction is out of these expressions, select the closest expression to it (among the 7 aforementioned expressions). If you cannot determine any expression, randomly select one of the mentioned expressions. If the predicted expression is another word in the family of the aforementioned expressions (happy, sad, surprise, fear, disgust, anger, neutral) change it to the corresponding word. For example, fearful is equal to fear, disgusted is equal to disgust, angry is equal to anger, surprised is equal to surprise, scared is equal to fear, and so on. If you randomly select an expression, do not need to explain, only generate the expression name.
@@ 4.2.2 Facial Expression Recognition
Figure 8: Confusion matix of each model across all datasets (in percent). In this experiment, all the datasets were used for both of the training and validation stages. No model was able to achieve high scores for the two challenging expressions, Fear and Disgust. A comparison of the results reveals that GPT-4o-mini and ViT models demonstrated the most consistent accuracy scores, while CLIP struggled to achieve satisfactory performance.
The average accuracy for each dataset, represented by $\mu$ in Fig. 7, demonstrates a correlation between dataset richness and complexity. Richer datasets tend to be more challenging, leading to lower average accuracies. The last two columns of this figure, labeled Mean and STD, summarize the mean accuracy and standard deviation for each model. Through this figure, POSTER and CEPrompt (two FER-specialized models) achieved the highest accuracy, which looks normal. Among the rest of 6 basic models, GPT-4o-mini achieved the highest accuracy, with minimum standard deviation. While this suggests POSTER, CEPrompt and GPT-4o-mini may be the fairest models, subsequent experiments reveal they exhibit significant bias.
In contrast, the CLIP model showed the worst performance, with an average accuracy of 49.7% and a high standard deviation exceeding 25%. These results indicate that the CLIP model is not well-optimized for FER tasks in its current form. Training the entire model, rather than just fine-tuning its decision-making head, could potentially improve its performance. Lastly, among the fully trained models (MobileNet, ResNet, XceptionNet, and ViT), ViT achieved the highest accuracy score, while ResNet exhibited the lowest standard deviation. Based solely on high accuracy and low standard deviation metrics, the models can be ranked from least to most accurate models as follows: POSTER, CEPrompt, GPT-4o-mini, ViT, ResNet, MobileNet, XceptionNet, and CLIP. We will discuss this ranking in Sec. 4.2.4 when we extract the metric-based bias scores.
TABLE IV: General accuracy of each model (in percent): four models—MobileNet, ResNet, XceptionNet, and ViT—were fully trained; three models—CLIP, POSTER, and CEPrompt—were fine-tuned; and GPT-4o-mini was evaluated solely on the validation sets without fine-tuning.
@@
POSTER 79.9 89.1 62.9 64.4 48.2 29.1 63.5 62.4 18.2
CEPrompt 79.3 88.8 61.7 68.4 42.4 29.2 62.7 61.7 18.9
TABLE V: Bias analysis using Equalized Odds, in the range of [0, 100], across models and three demographic attributes: age, gender, and race. Bias is calculated for each expression, followed by the calculation of maximum value, mean, and standard deviation for each attribute.
@@
Age 6.4 1.1 0.4 0.5 2.1 0.4 2.8 6.4 1.9 2.0
Gender 5.9 10.6 1.2 2.0 1.5 0.4 6.6 10.6 4.0 3.4
POSTER Race 10.8 6.9 6.9 4.8 5.6 2.1 6.3 10.8 6.2 2.4
Age 5.1 1.0 0.1 0.6 2.0 0.3 3.0 5.1 1.7 1.6
Gender 7.3 10.2 1.7 1.8 1.6 0.2 6.9 10.2 4.2 3.5
CEPrompt Race 12.4 6.3 6.1 5.4 4.4 2.1 6.4 12.4 6.1 2.9
Deeper analysis of Fig. 8 reveals significant confusion among certain expressions. For instance, pairs like Sad-Neutral, Disgust-Neutral, and Disgust-Anger pose considerable challenges, with erroneous predictions in the range of $18.8\%-35.3\%$, while zero values would be expected. The expressions accuracy below 50% are highlighted by red color in this figure. A comparison between POSTER and CEPrompt models and the rest of the models illustrates that they could not tackle the problem of Fear and Disgust expressions in FER task. On the other hand, while Fig. 7 showed, excluding two FER-specialized models, GPT-4o-mini as the top performer across various datasets, the general model results, summarized in Table IV, indicate that the ViT model achieved the highest average accuracy in this experiment. Additionally, MobileNet had the lowest standard deviation, indicating more consistency in its predictions.
TABLE VI: Bias analysis using Equal Opportunity, in the range of [0, 100], across models and three demographic attributes: age, gender, and race. Bias is calculated for each expression, followed by the calculation of maximum value, mean, and standard deviation for each attribute.
@@
Age 5.6 0.5 0.4 5.0 1.9 1.3 0.2 5.6 2.1 2.0
Gender 1.6 7.8 1.4 8.9 10.2 6.0 11.6 11.6 6.7 3.7
POSTER Race 6.0 8.2 12.1 24.3 25.0 19.2 8.0 25.0 14.6 7.4
Age 2.8 0.3 2.8 7.9 0.4 5.6 0.4 7.9 2.8 2.7
Gender 5.3 6.8 1.7 5.9 8.5 7.2 11.6 11.6 6.7 2.8
CEPrompt Race 8.8 8.5 11.0 24.2 20.0 9.5 6.6 24.2 12.6 6.1
So far, we have assessed the models’ accuracy to evaluate their robustness against bias stemming from imbalanced data distribution. Our findings indicate that, overall, POSTER and CEPrompt are the most accurate models, while CLIP performed poorly. Although these results suggest that POSTER, CEPrompt, GPT-4o-mini, and ViT might be the least biased models, subsequent experiments reveal different outcomes. In the next step, we will apply various bias metrics to further analyze the models’ biases across different attributes.
4.2.3 Metric-Based Model Analysis
Using the expression results from the previous experiments and the age, gender, and race attributes extracted with DeepFace [63], we evaluated the bias of all eight models with the metrics introduced in Sec. 2.2.
The first metric, Equalized Odds, examines the fairness of True Positive (TP) and False Positive (FP) rates across different groups. Table V highlights that Neutral and Happy were the most biased expressions. Almost all maximum biases for demographic attributes originate from these two expressions. For the Age attribute, Surprise emerged as the least biased expression, whereas for Gender and Race attributes, Fear showed the lowest bias. The Max column of this table reveals that the Age attribute consistently exhibited the lowest bias across all models. The Mean and STD columns further emphasize that, on average, Race was the most biased attribute across all models, while Age demonstrated the least bias. When comparing the Mean and STD values among the models, the Equalized Odds metric indicates that the ViT model was the most biased, while ResNet and XceptionNet demonstrated the lowest bias scores.
TABLE VII: Bias analysis using Demographic Parity, in the range of [0, 100], across models and three demographic attributes: age, gender, and race. Bias is calculated for each expression, followed by the calculation of maximum value, mean, and standard deviation for each attribute.
@@ 4.2.3 Metric-Based Model Analysis
Age 6.0 0.7 0.6 0.7 2.0 0.3 3.1 6.0 1.9 1.9
Gender 8.9 14.1 1.8 2.7 1.2 0.4 6.9 14.1 5.1 4.6
POSTER Race 12.3 7.2 5.5 5.4 5.3 2.0 4.5 12.3 6.0 2.9
Age 4.9 0.7 0.2 0.9 1.8 0.2 3.2 4.9 1.7 1.6
Gender 9.8 13.7 0.3 2.4 1.3 0.2 7.0 13.7 4.9 4.9
CEPrompt Race 13.2 7.0 5.8 5.8 4.4 1.6 5.7 13.2 6.2 3.2
Another metric we introduced in Sec.3.2 is Equal Opportunity, which focuses on the ratio of TP predictions across different groups. Information about this metric is provided in Table VI. Unlike Equalized Odds, this table shows that the Happy expression was not the most biased in terms of Equal Opportunity. For the Age attribute, Neutral ranked highest in bias, while for Gender and Race attributes, Happy and Surprise expressions exhibited the highest bias scores, respectively. Examining the Max, Mean, and STD columns of this table reveals that Race is the most bias-affected attribute across all models. Based on this metric, GPT-4o-mini exhibited the highest bias, while ResNet and MobileNet achieved the lowest bias scores.
Demographic Parity analyzes the likelihood of any group receiving a particular outcome. The data in Table VII shows that, compared to other bias metrics, Demographic Parity results in lower bias scores. For the Age attribute, the highest bias was observed in the Neutral expression, while Happy, Sad, and Surprise had minimal effects. In contrast, Happy was the most biased expression for the Gender attribute, while Fear and Sad showed the least bias. For the Race attribute, Neutral and Happy were the primary sources of bias, while Fear exhibited the lowest bias across all models. Although the Max column highlights higher biases for the Gender attribute, the Mean column underscores Race as the most biased attribute. As with the other bias metrics, Age remained the least biased attribute. The Mean and STD columns illustrate that for the Age attribute, the CLIP model was the most biased model, while for Gender and Race, the ViT model was the most biased model. Overall, considering the Mean bias scores, the ViT model emerged as the most biased, while ResNet demonstrated the least bias.
TABLE VIII: Bias analysis using the Treatment Equality, in the range of [0, 100], across models and three demographic attributes: age, gender, and race. Bias is calculated for each expression, followed by the calculation of maximum value, mean, and standard deviation for each attribute.
@@
Age 7.4 5.9 12.2 1.0 13.3 6.1 5.2 13.3 7.3 3.9
Gender 1.4 31.9 20.5 6.5 6.4 0.5 14.3 31.9 11.6 10.5
POSTER Race 10.7 37.1 24.3 30.7 37.5 12.2 39.4 39.4 27.4 11.1
Age 1.6 6.4 6.1 4.9 12.7 4.6 3.4 12.7 5.6 3.2
Gender 8.7 23.1 6.3 17.8 8.3 8.6 21.7 23.1 13.5 6.5
CEPrompt Race 14.4 37.3 20.9 46.0 37.3 16.5 34.8 46.0 29.6 11.2
The final bias measure we examined is Treatment Equality, which assesses the distribution of misclassified samples across different groups. Table VIII demonstrates that Treatment Equality resulted in higher bias scores. A detailed expression-level analysis shows that Sad, Fear, and Disgust were the main sources of bias for the Age attribute. Happy exhibited the highest bias for the Gender and Race attributes.Among the attributes, Race had the highest bias scores, while Age exhibited the lowest. This trend was consistent with our findings in the Equalized Odds and Equal Opportunity metrics. Comparing the Max column in Table VIII with the Max columns in Tables VVI, and VII shows that Treatment Equality exhibited higher bias scores across all attributes. In a model-wise comparison of Treatment Equality, CLIP showed the lowest bias score, while the highest bias was observed in ViT.
TABLE IX: Bias score for all models using the measures Equalized Odds (Eq-Od), Equal Opportunity (Eq-Op), Demographic Parity (De-Pa), and Treatment Equality (Tr-Eq) across age, race, and gender attributes. The mean represents the average of each row, and the bias score is the final mean score of the attributes. Data are in the range of [0, 100].
@@
Age 6.4 5.6 6.0 13.3 7.8
Gender 10.6 11.6 14.1 31.9 17.0
POSTER Race 10.8 25.0 12.3 39.4 21.8 15.5
Age 5.1 7.9 4.9 12.7 7.6
Gender 10.2 11.6 13.7 23.1 14.6
CEPrompt Race 12.4 24.2 13.2 46.0 23.9 15.4
To sum up, we analyzed bias of the models for each attribute and expression, followed by a model-based bias comparison. Based on our experiments, ViT showed the highest bias in the Equalized Odds, Demographic Parity, and Treatment Equality metrics, while GPT-4o-mini exhibited significant bias in Equal Opportunity. On the other hand, ResNet was the most robust model against bias in three out of four bias metrics.
Investigating bias scores for FER-specialized models, POSTER [31] and CEPrompt [32], reveals that despite achieving high accuracies in FER task, they suffer bias problem. Base on the ranking of models in each bias metric they mainly are located between the third to sixth place out of 8 models, which is not considerable. A peer-to-peer comparison between POSTER and CEPrompt showed that in three out of four metrics CEPrompt showed lower bias than POSTER.
Table IX presents the final bias score of each model across different attributes and bias metrics. Based on the Eq. 12, we used the maximum bias score of each model as its bias score. In the next step, using the Eq. 13, we calculated the average bias score for each model across all attributes. Finally, through the Eq.  14, the overall bias score for each model was obtained by averaging the bias scores for each attribute. The final bias scores showed that GPT-4o-mini was the most biased model, followed by ViT in the second place. Additionally, this table highlights that residual-based models, including ResNet and XceptionNet, showed the minimum bias. The final bias score for the POSTER and CEPrompt is not considerable, where they ranked as the fourth and third least bias models. Ultimately, our experiments rank the models from the most biased to the least biased as follows: GPT-4o-mini, ViT, CLIP, MobileNet, POSTER, CEPrompt, XceptionNet, and ResNet. Notably, this research focused solely on the FER task, and the observed bias in the models could be evaluated on different tasks, potentially yielding varying bias scores.
4.2.4 Conclusion
The analysis of the eight FER models demonstrates that all exhibit varying degrees of bias across Age, Gender, Race, and expression attributes. While specialized models such as POSTER and CEPrompt achieve higher accuracy, they do not eliminate bias. Overall, GPT-4o-mini and ViT were the most biased models, whereas ResNet and XceptionNet showed the least bias. These results highlight that high accuracy does not necessarily correspond to fairness, emphasizing the need to evaluate both performance and bias when assessing FER models.